# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

### v3

@@ 4.3 Scenario Modeling
Next, we define the state space, action space, and return function, thereby modeling the driving scenario as a Markov Decision Process (MDP).
@@ 4.3.1 State
The state $s_{t}$ encompasses observations of the ego vehicle, including driving information features from both the ego vehicle and surrounding vehicles. As described in Section 4.2, at time $t$, $s_{t}$ is defined as follows:
@@
where $v_{e}$ is the speed of the ego vehicle. The speeds of surrounding vehicles and their longitudinal positions relative to the ego vehicle are represented as $\{v_{i},\ d_{i}\}$, where $i=(1,2,3,4,5)$ corresponds to the front vehicle, left front vehicle, left rear vehicle, right front vehicle, and right rear vehicle, respectively. To obtain more accurate state features, we make the following assumptions:
  1. 1.
    When the ego vehicle is driving in the nearside lane, one side of the ego vehicle will have no surrounding vehicles. In this case, $\{v_{i},\ d_{i}\}$ for vehicles on that side will be set to $0\ \text{m}\}$/s , 100 m }.
  2. 2.
    If the ego vehicle is not driving in the nearside lane and there are no other vehicles in a side lane, or if vehicles in a side lane are outside of the perception range, the $\{v_{i},\ d_{i}\}$ for vehicles on that side will be set to $0\ \text{m}\}$/s , 50 m }.
@@ 4.3.2 Action
In this paper, the decision-making layer provides only discrete meta-actions, which serve as high-level commands for the path planning and control layer. Consequently, the high-level action $a_{t}$ of the ego vehicle at time $t$ can be defined as follows:
@@ 4.3.3 Return
In the highway lane-change scenario, the primary objective of autonomous vehicles is to achieve high-speed driving while avoiding collisions. Therefore, to guide the ego vehicle in increasing speed while maintaining a high level of safety, the return function in this paper is composed of two parts:
@@
    Efficiency Reward: The maximum speed of the ego vehicle is set to 25 m/s. To encourage the ego vehicle to accelerate from 0 m/s to higher speeds, we employ an efficiency reward function, defined as follows:
@@
    where $\alpha_{1}>0$ represents the weight coefficient used to adjust the importance of efficiency reward in the strategy.
@@
    Safety Cost: The safety cost depends on the safe distance $d_{safe}$, which represents the longitudinal distance between the ego vehicle and the vehicles in front or rear, as well as whether a collision occurs. The safety cost function can be defined as follows:
@@
    where $\alpha_{2}>0$ represents the weight coefficient used to adjust the importance of safety cost in the strategy.
The weight coefficients $\alpha_{1}$ and $\alpha_{2}$ are adjusted based on the results of multiple experiments, and their specific values are provided in Table 1. The final return function $R$ is presented in Eq. 32:
@@ 4.4 Simulation Settings
All experiments in this paper were conducted on a computer equipped with an Intel i5-13600KF CPU, an NVIDIA GeForce RTX 4090 GPU, and 32GB of RAM. The experiments involved two simulation environments, Highway-Env and Carla, with their respective runtime parameters listed in Table 5 and Table 5. As this paper focuses on decision-making methods, we simplified the perception module by assuming that the ego vehicle can directly obtain accurate information (e.g., speed, distance) about other vehicles within a 50 m radius. To demonstrate the algorithm’s generalization ability, different random seeds were employed each time the environment was reset. In the Carla simulation environment, all vehicles except the ego vehicle were managed by the simulator’s traffic manager, which controls behaviors such as car-following and lane-change. In the Highway-Env simulation environment, other vehicles were managed using simple decision-making algorithms (e.g., IDM, MOBIL) and control algorithms (e.g., proportional controllers).
@@
Auto lane change for other vehicle True
@@ 5 Simulation Result
TS2C (Teacher-Student Shared Control (Xue et al., 2023), TSF): TS2C is a student-teacher framework where a pre-trained teacher agent guides the student agent by intervening and providing online demonstrations based on trajectory-based value estimation, and we re-implemented this algorithm using PPO as the base method.
@@
  • 1)
    Model Training: Each algorithm was trained 3 times, using different random seeds to capture their average performance. The model is tested every 5,000 steps during the training process, and the results are recorded to evaluate its training efficiency and performance.
  • 2)
    Model Evaluation: Following training, each algorithm was evaluated 6 times with different random seeds to ensure an accurate assessment of average performance. These test results primarily evaluate the models’ success rates in safely driving a target distance of 1 km without collisions, while also assessing traffic efficiency and driving safety throughout the process. Additionally, all models were evaluated in both high-density and low-density highway scenarios.
@@ 5.1 Model Training
Figure 4: The training curves for all algorithms are presented. Each algorithm is trained with three different random seeds, accumulating a total of 500K training steps for each case, except for S2CD, which requires only 300K training steps. To ensure a comprehensive assessment of algorithm performance, two evaluation episodes are conducted for every 5,000 training steps. The average value of these two evaluation episodes is taken as the result.
@@
Throughout the training process of the algorithms, we primarily focus on the trends of key indicators, such as model performance and safety. Fig. 4(a) illustrates the test return curve as a function of training steps, clearly reflecting the overall performance of all algorithms. As the curve indicates, guided by the teacher model, S2CD benefits from higher sample efficiency and learning efficiency, enabling it to acquire more accurate knowledge in the early stages. This advantage allows it to cover longer distances without collisions, thereby achieving consistently high return values of approximately 240 from the outset. For the TS2C algorithm, excessive teacher intervention during the early stages leads to an imbalance between positive and negative samples, and the discrepancy between the policy distributions of the teacher and the student hampers the student’s ability to optimize its policy, resulting in a lower Return. SOAR-ACPPO decides whether to intervene based on intervention probability, which prevents it from strictly following the teacher’s guidance, consequently leading to a lower initial score. Nevertheless, SOAR-ACPPO’s learning speed is significantly faster than that of other algorithms, except for S2CD. In contrast, other RL algorithms gradually optimize their policies through environmental interactions. However, due to limited early-stage knowledge, they experienced more frequent collisions during testing, resulting in significantly lower return values compared to S2CD. Furthermore, Fig. 4(d) demonstrates that the efficiency reward obtained by the S2CD and TS2C agents through interactions with the environment are considerably higher than those achieved by other algorithms, indicating that the teacher’s intervention effectively enhances the student’s ability to achieve high scores. Regarding safety, Fig. 4(e) illustrates the safety cost curves during the data collection period. Due to the teacher model’s intervention in correcting the student’s unsafe actions, S2CD and TS2C ensure that the ego vehicle maintains a safe distance from other vehicles whenever feasible. Consequently, this leads to their safety cost during the data collection period being significantly lower than that of other algorithms. However, as depicted in Fig. 4(b), the TS2C algorithm demonstrates poor performance in terms of test safety cost, indicating the model’s inefficiency in learning. In contrast, S2CD consistently achieves the lowest test safety cost. Fig. 4(f) illustrates that during the early data collection phase, S2CD experienced the least number of collision events. Additionally, TS2C and SOAR-ACPPO, both benefiting from teacher guidance, also exhibited a similarly low collision probability, approximately half that of other algorithms. Therefore, even in the initial stages of training, when the policy is not yet fully optimized, the teacher’s guidance enables S2CD’s student agent to achieve performance and safety significantly superior to those of other algorithms.
@@
During the mid-to-late stages of model training, as agents explore the environment more thoroughly, accumulate data, and update their policies, the performance of other algorithms, excluding DQN, begins to improve. Specifically, return, efficiency reward, and driving speed gradually increase, while safety cost and collision counts decrease. Notably, the PPO algorithm and its derivatives: S2CD, TS2C, SOAR-ACPPO, PPO-Lag, and GAIL, demonstrate significantly faster performance improvements, whereas SAC and its derivative SAC-Lag learn at a slower pace. This is because PPO-based algorithms limit the magnitude of policy updates, preventing substantial performance degradation in each iteration and ensuring that the policies steadily progress toward better solutions until convergence. Among these, TSF-based algorithms (S2CD, TS2C, SOAR-ACPPO) demonstrated relatively superior final performance. In contrast, because SAC-based algorithms are inherently designed for continuous action tasks, achieving satisfactory training results for the discrete action tasks in this study is challenging. In terms of training efficiency, S2CD’s enhancements in sampling and learning mechanisms result in a noticeably faster convergence speed, achieving a return greater than 290 with only 300K steps. This sharply contrasts with other algorithms, which require 500K steps to achieve convergence while exhibiting inferior performance. Although the safety cost of S2CD experiences a slight increase during the intermediate stages, it subsequently declines in the later stages and consistently remains lower than that of other algorithms. Notably, S2CD experiences significantly fewer collisions than other algorithms during the entire training process, with only approximately 2 collisions occurring every 5,000 training steps in the later stage. This indicates that S2CD demonstrates exceptionally high safety performance during the training process, markedly outperforming other algorithms.
@@ 5.2 Model Evaluation
After completing the training, we conducted a final performance evaluation of all models, as illustrated in Fig. 5 and presented in Table 6. It is worth noting that the offline RL algorithm CQL was trained using data collected online by the SAC algorithm. However, due to the poor performance of the SAC algorithm, CQL generated actions that consistently remained in "follow" mode even after training, failing to perform any lane-change actions. Therefore, we exclude its performance from further comparisons.
@@
To validate the generalizability and superiority of the S2CD framework, we evaluated the performance of all compared algorithms in both high-density and low-density traffic scenarios. In the high-density traffic scenario, vehicle spacing was randomly set within a range of 20 m to 50 m, resulting in an average vehicle density of approximately 29 vehicles per kilometer per lane. In contrast, in the low-density traffic scenario, vehicle spacing was randomly set within a range of 90 m to 120 m, resulting in an average vehicle density of approximately 9 vehicles per kilometer per lane. The average evaluation results for these scenarios are presented in s Table 7 and Table 8, respectively. As shown in Table 7, in the high-density traffic scenario, the reduced vehicle spacing increases the complexity of the environment, further degrading the performance of the already underperforming DQN, SAC-Lag and GAIL algorithms. However, as lane-change opportunities become more limited, the frequency of lane-change by the ego vehicle significantly dropped, resulting in a reduced probability of collisions and, consequently, an improvement in the performance of other algorithms. S2CD continues to demonstrate the highest success rate (96.83%) and return value (143.88), while its safety cost (0.77) and speed (15.02 m/s) are also very close to the best values of 0.63 and 15.11 m/s. As shown in Table 8, in the low-density traffic scenario, the sparsity of traffic flow reduces the necessity for frequent lane-change to increase speed, thereby decreasing the probability of collisions and improving the safety performance and driving speed of all algorithms. S2CD achieved the highest success rate (98.75%) and return value (341.01), while also demonstrating the lowest safety cost (0.51). Additionally, its driving speed (21.32 m/s) is close to the highest speed achieved by the PPO algorithm, which is 21.68 m/s. Similarly, the overall performance of TSF-based TS2C and SOAR-ACPPO ranks second only to S2CD, outperforming other algorithms in both high-density and low-density traffic scenarios. This finding aligns with the results in the medium-density traffic scenario, further validating the effectiveness of teacher guidance.
@@ 5.3 Ablation Study
To assess the contributions of each component in the S2CD framework and validate its effectiveness, we conducted ablation experiments. By gradually removing or replacing specific modules within the framework, we observed variations in performance. The ablation study validated the framework’s design rationale and demonstrated the role that each module plays in enhancing performance.
@@ 5.3.1 The Impact of Different Teacher Quality
Table 9 presents the computational costs associated with different teacher models, including training steps, number of processes, time spent, GPU memory footprint, memory footprint, and CPU utilization. In the Highway-Env environment, training a low-quality teacher model requires 200K steps, whereas achieving a high-quality teacher model can be accomplished by simply doubling the number of data collection processes. Although the training steps and CPU utilization approximately doubled, the increases in time spent, GPU memory footprint, and memory footprint were minimal, with the training time remaining almost unchanged. Therefore, in the Highway-Env environment, a high-quality teacher model can be obtained at minimal computational cost. In contrast, training for the same 400K steps in the Carla simulation environment required nearly 5 times the time spent compared to Highway-Env, with the GPU memory footprint, memory footprint, and CPU utilization more than doubling. This indicates that the cost of training teacher models in complex simulation environments is significantly greater than in simpler ones. The 3 trained teacher models were integrated into the S2CD framework to guide the student agent’s learning in the Carla environment, with the training and evaluation results presented in Fig. 6 and Table 10.
@@
Figure 6: The training and evaluation processes involve different teacher models. 1. Model Training: Each algorithm is trained with three different random seeds, accumulating a total of 300K training steps for each case. Two evaluation episodes are conducted for every 5,000 training steps. The average value of these two evaluation episodes is taken as the result. During model training, we record the return value, safety cost, and collision counts. 2. Model Evaluation: Each model trained with varying random seeds is evaluated twice. Each algorithm undergoes six evaluation runs, with each evaluation consisting of 200 episodes. For model evaluation, we record the success rate.
@@
The results demonstrate that higher-quality teacher models enable the student to acquire more accurate knowledge, resulting in improved performance during the later stages of training and in the final evaluation. Specifically, the return, the probability of avoiding collisions and the task success rate during training were higher with the high-quality teacher compared to the low-quality teacher. This indicates that high-performance teachers can significantly enhance the safety and performance of the student’s training. Additionally, we implemented a weaning mechanism that gradually reduces the teacher’s influence on the student agent in the later stages, allowing the student to explore the environment more independently. This mechanism elevated the algorithm’s performance ceiling, enabling the student to ultimately surpass the teacher in key metrics such as return and success rates. Notably, although the teacher model trained in the Carla environment consumed significant computational resources, its performance was inferior to that of the low-quality teacher trained in the Highway-Env environment. The return and success rates were only 238.64 and 66.00%, respectively, significantly lower than those of the teachers trained in simpler environments. Furthermore, when guided by this teacher model, the student agent’s final performance was lower than that of the simple teacher. The return was approximately 40 lower, and the key metric, success rate, was about 20% lower. This demonstrates that training teacher models in a simple simulation environment not only saves substantial computational costs but also yields better performance, providing more effective guidance to the student agent.

### v4

@@ 4.3 Scenario Modeling
Next, we define the state space, action space, and return function, thereby modeling the driving scenario as a MDP.
@@ 4.3.1 State
The state $s_{t}$ includes information of vehicles. As described in Section 4.2, $s_{t}$ is as follows:
@@
where $v_{e}$ is the velocity of the ego vehicle; $\{v_{i},\ d_{i}\}$ represents the velocities and distances of surrounding vehicles, where $i=(1,2,3,4,5)$ corresponds to the front, front-left, rear-left, front-right, and rear-right vehicles, respectively.
@@ 4.3.2 Action
In this paper, the decision-making layer provides discrete commands for the low-level execution layer. Consequently, the action $a_{t}$ is defined as follows:
@@ 4.3.3 Return
In the highway lane-change scenario, the main goal of autonomous vehicles is to drive safely and at high speed. Therefore, the return function is designed with the following two components:
@@
    Efficiency Reward: The maximum velocity for all vehicles is capped at 25 m/s. To prompt the ego vehicle to achieve greater speeds, we adopted the following efficiency reward function:
@@
    where $\alpha_{1}>0$ is a hyperparameter that adjusts the significance of efficiency within the policy.
@@
    Safety Cost: To maintain a safe distance between the ego vehicle with surrounding vehicles and to prevent collisions, we define the safety cost function as follows:
@@
    where $\alpha_{2}>0$ is a hyperparameter that adjusts the significance of safety within the policy; $d_{safe}$ represents the safe distance, defined as the distance between the ego vehicle and the vehicle directly in front or behind.
The values of $\alpha_{1}$ and $\alpha_{2}$ are provided in Table 1, and the final Return function $R$ is presented in Eq. 29:
@@ 4.4 Simulation Settings
All simulation experiments were conducted on a computer equipped with an Intel i5-13600KF CPU, an NVIDIA GeForce RTX 4090 GPU, and 32GB of RAM. The experiments involved two simulation environments, Highway-Env and Carla, with their respective runtime parameters listed in Table 5 and Table 5. Given the focus of this paper on decision-making approaches, we streamlined the perception module by assuming the ego vehicle has direct access to accurate information (e.g., speed, distance) about surrounding vehicles within a 50-meter range. To demonstrate the algorithm’s generalization ability, different random seeds were employed each time the environment was reset. In the Carla environment, the car-following and lane-changing behaviors of other vehicles are managed by the Traffic Manager. In the Highway-Env simulation environment, other vehicles were managed using simple decision-making algorithms (e.g., IDM, MOBIL) and control algorithms (e.g., proportional controllers).
@@
Auto lane change True
@@ 5 Simulation Result
TS2C (Teacher-Student Shared Control (Xue et al., 2023), TSF): TS2C is a student-teacher framework where a pre-trained teacher agent guides the student agent by intervening and providing online demonstrations based on trajectory-based values, and we re-implemented this algorithm using PPO as the base method.
@@ 5.1 Model Training
Figure 3: The training curves for all algorithms are shown. Every algorithm is trained using three random seeds, accumulating 500K training steps in every case, except for S2CD, which needs only 300K steps. To evaluate the performance of the algorithms at each stage, two evaluation episodes are conducted every 5,000 training steps, and the average value of these episodes is recorded as the result.
@@
Throughout the training process of the algorithms, we primarily focus on the trends of key indicators, such as model performance and safety. Fig. 3(a) illustrates the test return curve as a function of training steps, clearly reflecting the overall performance of all algorithms. As the curve indicates, guided by the teacher model, S2CD benefits from higher sample efficiency and learning efficiency, enabling it to acquire more accurate knowledge in the early stages. This advantage allows it to cover longer distances without collisions, thereby achieving consistently high return values of approximately 240 from the outset. For the TS2C algorithm, excessive teacher intervention during the early stages leads to an imbalance between positive and negative samples, and the discrepancy between the policy distributions of the teacher and the student hampers the student’s ability to optimize its policy, resulting in a lower Return. SOAR-ACPPO decides whether to intervene based on intervention probability, which prevents it from strictly following the teacher’s guidance, consequently leading to a lower initial score. Nevertheless, SOAR-ACPPO’s learning speed is significantly faster than that of other algorithms, except for S2CD. In contrast, other RL algorithms gradually optimize their policies through environmental interactions. However, due to limited early-stage knowledge, they experienced more frequent collisions during testing, resulting in significantly lower return values compared to S2CD. Furthermore, Fig. 3(d) demonstrates that the efficiency reward obtained by the S2CD and TS2C agents through interactions with the environment are considerably higher than those achieved by other algorithms, indicating that the teacher’s intervention effectively enhances the student’s ability to achieve high scores. Regarding safety, Fig. 3(e) illustrates the safety cost curves during the data collection period. Due to the teacher model’s intervention in correcting the student’s unsafe actions, S2CD and TS2C guarantee that the ego vehicle keeps a safe distance from other vehicles whenever feasible. Consequently, this leads to their safety cost during the data collection period being significantly lower than that of other algorithms. However, as depicted in Fig. 3(b), the TS2C algorithm demonstrates poor performance in terms of test safety cost, indicating the model’s inefficiency in learning. In contrast, S2CD consistently achieves the lowest test safety cost. Fig. 3(f) illustrates that during the early data collection phase, S2CD experienced the least number of collision events. Additionally, TS2C and SOAR-ACPPO, both benefiting from teacher guidance, also exhibited a similarly low collision probability, approximately half that of other algorithms. Therefore, even in the initial stages of training, when the policy is not yet fully optimized, the teacher’s guidance enables S2CD’s student agent to achieve performance and safety significantly superior to those of other algorithms.
@@
During the mid-to-late stages of training, as agents accumulate data, and update their policies, the performance of other algorithms, excluding DQN, begins to improve. Specifically, return, efficiency reward, and driving speed gradually increase, while safety cost and collision counts decrease. Notably, the PPO algorithm and its derivatives: S2CD, TS2C, SOAR-ACPPO, PPO-Lag, and GAIL, demonstrate significantly faster performance improvements, whereas SAC and its derivative SAC-Lag learn at a slower pace. This is because PPO-based algorithms limit the magnitude of policy updates, preventing substantial performance degradation in each iteration and ensuring that the policies steadily progress toward better solutions until convergence. Among these, TSF-based algorithms (S2CD, TS2C, SOAR-ACPPO) demonstrated relatively superior final performance. In contrast, because SAC-based algorithms are inherently designed for continuous action tasks, achieving satisfactory training results for the discrete action tasks in this study is challenging. In terms of training efficiency, S2CD’s enhancements in sampling and learning mechanisms result in a noticeably faster convergence speed, achieving a return greater than 290 with only 300K steps. This sharply contrasts with other algorithms, which require 500K steps to achieve convergence while exhibiting inferior performance. Although the safety cost of S2CD experiences a slight increase during the intermediate stages, it subsequently declines in the later stages and consistently remains lower than that of other algorithms. Notably, S2CD experiences significantly fewer collisions than other algorithms during the entire training process, with only approximately 2 collisions occurring every 5,000 training steps in the later stage. This indicates that S2CD demonstrates exceptionally high safety performance during the training process, markedly outperforming other algorithms.
@@ 5.2 Model Evaluation
Upon completing model training, a final performance evaluation was conducted for all models, as illustrated in Fig. 4 and presented in Table 6. Notably, CQL consistently maintained the "follow" action during testing without performing any lane-change maneuvers. Therefore, it is excluded from subsequent analyses.
@@
To further validate the generalizability of the S2CD framework, we examined all algorithms in high-density and low-density scenarios. For high-density scenario, vehicle spacing ranged between 20 and 50 meters, with an average of about 29 vehicles per lane per kilometer. The evaluation results are shown in Table 7. In this scenario, the reduced vehicle spacing increases the complexity of the environment, further degrading the performance of the already underperforming DQN, SAC-Lag and GAIL algorithms. However, as lane-change opportunities become more limited, the frequency of lane-change by the ego vehicle significantly dropped, resulting in a reduced probability of collisions and, consequently, an improvement in the performance of other algorithms. S2CD continues to demonstrate the highest success rate (96.83%) and return value (143.88), while its safety cost (0.77) and speed (15.02 m/s) are also very close to the best values of 0.63 and 15.11 m/s. For the low-density scenario, vehicle spacing ranged from 90 to 120 meters, with an average of about 9 vehicles per lane per kilometer. The evaluation results are shown in Table 8. In this scenario, vehicles could achieve higher speeds without frequent lane changes, significantly reducing collision probability and thus improving the safety and efficiency of all algorithms. S2CD achieved the highest success rate (98.75%) and return value (341.01), while also demonstrating the lowest safety cost (0.51). Additionally, its driving speed (21.32 m/s) is close to the highest speed achieved by the PPO algorithm, which is 21.68 m/s. Similarly, TS2C and SOAR-ACPPO, which are based on the TSF framework, ranked just behind S2CD in overall performance, outperforming other algorithms. These results are consistent with those in the moderate-density scenario, further validating the effectiveness of teacher-guided learning.
@@ 5.3 Ablation Study
To assess the contributions of every part within the S2CD framework and validate its effectiveness, we conducted ablation experiments. By gradually removing or replacing specific modules within the framework, we observed variations in performance. The ablation study validated the framework’s design rationale and demonstrated the role that each module plays in enhancing performance.
@@ 5.3.1 The Impact of Different Teacher Quality
Table 9 presents the computational costs associated with different teacher models, including training steps, number of processes, time spent, GPU memory footprint, memory footprint, and CPU utilization. In the Highway-Env environment, training a low-quality teacher model requires 200K steps, whereas achieving a high-quality teacher model can be accomplished by simply doubling the number of data collection processes. Although the training steps and CPU utilization approximately doubled, the increases in time spent, GPU memory footprint, and memory footprint were minimal, with the training time remaining almost unchanged. Therefore, in the Highway-Env environment, a high-quality teacher model can be obtained at minimal computational cost. In contrast, training for the same 400K steps in the Carla simulation environment required nearly 5 times the time spent compared to Highway-Env, with the GPU memory footprint, memory footprint, and CPU utilization more than doubling. This indicates that the cost of training teacher models in complex simulation environments is significantly greater than in simpler ones. The three trained teacher models were integrated into the S2CD framework to guide the learning of the student in Carla, with the training and evaluation results presented in Fig. 5 and Table 10.
@@
Figure 5: The training and evaluation processes involve different teacher models: 1. Model Training: Every algorithm is trained with three random seeds, accumulating 300K training steps per case. Two evaluation episodes are conducted every 5,000 training steps, and their average value is recorded as the result. Throughout model training, we monitor the return value, safety cost, and collision counts. 2. Model Evaluation: Every model trained with different random seeds is evaluated twice. In total, every algorithm undergoes six evaluation runs, with each run consisting of 200 episodes. During model evaluation, the success rate is recorded.
@@
The results demonstrate that higher-quality teacher models enable the student to acquire more accurate knowledge, resulting in improved performance during the later stages of training and in the final evaluation. Specifically, the return, the probability of avoiding collisions and the task success rate during training were higher with the high-quality teacher compared to the low-quality teacher. This indicates that high-performance teachers can significantly enhance the safety and performance of the student’s training. Additionally, we implemented the weaning mechanism to gradually decrease the teacher’s influence on the student agent in the later stages, allowing the student to explore the environment more independently. This mechanism elevated the algorithm’s performance ceiling, enabling the student to ultimately surpass the teacher in key metrics such as return and success rates. Notably, although the teacher model trained in the Carla environment consumed significant computational resources, its performance was inferior to that of the low-quality teacher trained in the Highway-Env environment. The return and success rates were only 238.64 and 66.00%, respectively, significantly lower than those of the teachers trained in simpler environments. Furthermore, when guided by this teacher model, the student agent’s final performance was lower than that of the simple teacher. The return was approximately 40 lower, and the key metric, success rate, was about 20% lower. This demonstrates that training teacher models in a simple simulation environment not only saves substantial computational costs but also yields better performance, providing more effective guidance to the student agent.